# CUORE Collaboration

### v1

@@ VIII Low Energy Spectra
The 90 keV events likely originate from nuclear recoil from 210Po $\alpha$-decay on crystal surfaces. This explanation gains support from the fact that the peak structure is more pronounced in the M2 spectrum compared to the M1, suggesting a surface contamination. The hypothesis has been verified by looking at the energy of M2 coincident events, which peaks at $\sim$5.3 MeV, the energy of the $\alpha$ emitted by the 210Po.
@@
The monochromatic peak at $\sim$31 keV was observed as well in CUORE-0 demonstrators and past CUORE crystals measurements [9, 37] and ascribed to a physics feature. Specifically, in Ref. [37] it was found that a few events were in coincidence with high energy $\gamma$s (507.6 keV and 573.1 keV) associated with 121Te electron capture (EC), suggesting it to correspond to the binding energy of the K-shell of 121Sb, consequent to a K-shell EC. Moreover, in Ref. [41], a peak at the same energy was used to claim the discovery of 123Te EC, later refuted in Ref. [40] and ascribed to the one of 121Te, activated by neutrons. Its observation in CUORE indicates that the origin of this peak is still present in either TeO2 crystals or the detector setup. We observe its rate to be approximately constant along the data taking with a weighted average among the datasets of 1.35 $\pm$ 0.04 counts/(kg d), while we estimate its mean to be 30.88 $\pm$ 0.09 keV. The reconstruction of this peak is affected by its proximity to the broad structure around $\sim$36 keV. As a systematic test, we repeat the fit by narrowing the energy range and employing the stricter selection with a 3 keV threshold, to take advantage of the improved resolution. Both tests report a $\sim$30lower rate, and a $\sim$0.8lower value of the mean, that we account as a systematic uncertainty.
Focusing on the detectors available down to 3 keV, as in Fig. 14, thanks to a stronger background suppression and improved energy resolution, we observe two excess structures at 10 and 13 keV. A similar feature, hidden by the background rise, can be found in the CUORE-0 spectrum in Ref. [9]. We model these features using two Gaussian distributions, fixing the resolution to the one of the baseline, and a linear background. Their means are reconstructed to be (10.3 $\pm$ 0.2) and (12.6 $\pm$ 0.3) keV, with a weighted average rate of (0.66 $\pm$ 0.07) and (0.67 $\pm$ 0.07) counts/(kg d).
We observe a peak on the rising edge close to 3 keV. We modeled it as a Gaussian with resolution fixed to that of the baseline and overlaid it with a linear and an exponential component as the background. We reconstruct a mean and a rate, as the weighted average among the datasets, of 4.69 $\pm$ 0.03 keV and 19.1 $\pm$ 1.0 counts/(kg d), respectively. The same peak was observed in past CUORE crystals measurements and the demonstrator Cuoricino [37]. Similar to the peak at 30.5 keV, it was ascribed to the L1-shell binding energy of 121Sb, characteristic to the L1-shell EC of 121Te. It was also supported by the fact that few events were found to be coincident with high energy $\gamma$s associated to this process.
We look for coincidences with 121Te EC $\gamma$ peaks in CUORE data and we find overall 20 events, of which 4 between 30 and 32 keV, 1 between 3 and 6 keV, and the rest spread up to 40 keV. The low significance of this result makes any conclusions drawn from this study unreliable, leaving open the possibility of an alternative explanation.
@@ IX Conclusion and Outlook
We studied the main spectral components, which will be fundamental for further investigations about rare nuclear decays (e.g. 123Te electron capture), new physics searches (e.g. solar axions [43, 44], WIMPs [32, 45] etc.), and more generally about keV-scale background predictions for future tonne-scale cryogenic calorimeter experiments, such as CUPID ([46, 47]). This result enables the development of a detailed background model inspired by the work in Ref. [16] for energies lower than 100 keV.
@@
CUORE will stop data-taking after reaching 3 tonne yr of TeO2 exposure to undergo major upgrades to the cryogenic facility, mainly devoted to the pulse tubes pre-cooling system. This is motivated by the need of its successor, CUPID [18, 19], for a lower vibration noise to allow the successful operation of light detectors [48, 49, 50]. Such an upgrade should be particularly beneficial for low-energy studies in CUORE, serving as a proof of principle for future large-mass cryogenic calorimeters operating in similar facilities. In particular, the noise suppression offered by the cryostat refurbishment can increase the number of available detectors at this energy scale, and thus the sensitivity to beyond Standard Model physics. The background studies can also profit from the improved multi-site event-tagging capability that will be made possible by a higher number of active detectors at lower thresholds.
A second phase of CUORE, after the cryogenic system upgrade, can pursue low energy studies with an improved noise environment. The results presented in this work provide key insights for operating the detectors to maximize the performance for keV-scale physics, together with data analysis methods and strategies. The tools we developed will be used to understand the potential for CUORE and future tonne-scale cryogenic calorimeters as multipurpose experiments, demonstrating the scalability of this particle detector’s technology in view of both rare decays and dark matter interactions.

### v2

@@ VI Low Energy Spectra
The $\sim$90 keV events likely originate from nuclear recoil from 210Po $\alpha$-decay on crystal surfaces. This explanation gains support from the fact that the peak structure is more pronounced in the M2 spectrum compared to the M1, suggesting a surface contamination. The hypothesis has been verified by looking at the energy of M2 coincident events, which peaks at $\sim$5.3 MeV, the energy of the $\alpha$ emitted by the 210Po.
@@
[…]
Focusing on the detectors available down to 3 keV, as in Fig. 14, thanks to a stronger background suppression and improved energy resolution, we observe two excess structures at $\sim$10 and $\sim$13 keV. A similar feature, hidden by the background rise, can be found in the CUORE-0 spectrum in Ref. [8]. We model these features using two Gaussian distributions, fixing the resolution to the one of the baseline, and a linear background. Their means are reconstructed to be (10.3 $\pm$ 0.2) and (12.6 $\pm$ 0.3) keV, with a weighted average rate of (0.66 $\pm$ 0.07) and (0.67 $\pm$ 0.07) counts/(kg d).
We observe a peak on the rising edge close to 3 keV. We modeled it as a Gaussian with resolution fixed to that of the baseline and overlaid it with a linear and an exponential component as the background. We reconstruct a mean and a rate, as the weighted average among the datasets, of 4.69 $\pm$ 0.03 keV and 19.1 $\pm$ 1.0 counts/(kg d), respectively.
The same peak was observed in past CUORE crystals measurements and the demonstrator Cuoricino. Similar to the peak at $\sim$31 keV, in [13], a few events in this peak were found to be coincident with higher energy gammas, consistent with the L1-shell EC of 121Te. However, the observed excess of events in this peak, not coincident with these gammas, was not explicitly attributed to 121Te EC.
We look for coincidences with the same 121Te EC $\gamma$ peaks in CUORE data and we find overall 20 events, of which 4 between 30 and 32 keV, 1 between 3 and 6 keV, and the rest spread up to 40 keV. The low significance of this result makes any conclusions drawn from this study unreliable, leaving open the possibility of an alternative explanation.
@@ VII Conclusion and Outlook
We studied the main spectral components, which will be fundamental for further investigations about rare nuclear decays (e.g. the aforementioned 123Te electron capture) and new physics searches, briefly described in the following.
A search for Weakly Interacting Massive Particle (WIMP) dark matter can be performed by looking for the annual modulation signature caused by the seasonal variation in the velocity distribution of WIMP interactions with the detector nuclei due to the Earth’s revolution around the sun [33]. The WIMP scattering interactions will create observable nuclear recoils, both on Te and O nuclei, below $\sim$50 keV. Therefore a low threshold is necessary for an increased sensitivity to these scattering interactions [42, 30] . CUORE can leverage its 5 years long data taking to observe multiple cycles of this phenomenon.
Another promising dark matter candidate is the axion, a pseudoscalar boson originally proposed to solve the strong CP problem [35]. This particle can be produced by nuclear reactions in the core of Sun and propagate to the Earth, eventually converting into an electron. The most intense signature is foreseen from the de-excitation of thermally populated first excited state of 57Fe, which would produce a 14.4 keV monochromatic peak in CUORE crystals.
Moreover, non-relativistic axions might compose the dark matter halo in the galaxy and interact with matter releasing an amount of energy corresponding to their mass. Performing a scan over accessible energies, CUORE can search for dark matter axions at specific mass values.
[…]
More generally, this work enables the development of a detailed background model inspired by Ref. [6] for energies lower than 100 keV, which can be employed for new physics sensitivity studies with future tonne-scale cryogenic calorimeter experiments, such as CUPID ([24, 23]).
@@
Beyond paving the way to new physics investigation at the keV-scale with CUORE data, these results inform planning and goals for future stages of large mass cryogenic calorimeters. CUORE will stop data-taking after reaching 3 tonne yr of TeO2 exposure to undergo major upgrades to the cryogenic facility, mainly devoted to the pulse tubes pre-cooling system, that is needed to cool down about a tonne of material down to 4K. This is motivated by the need of its successor, CUPID [19, 18], for a lower vibration noise to allow the successful operation of light detectors [16, 15, 17]. Such an upgrade should be particularly beneficial for low-energy studies in CUORE, serving as a proof of principle for future large-mass cryogenic calorimeters operating in similar facilities. In particular, the noise suppression offered by the cryostat refurbishment can increase the number of available detectors at this energy scale, and thus the sensitivity to beyond Standard Model physics. The background studies can also profit from the improved multi-site event-tagging capability that will be made possible by a higher number of active detectors at lower thresholds.
The results presented in this work provide key insights for operating the detectors to maximize the performance for keV-scale physics, together with data analysis methods and strategies. The tools we developed will be used to understand the potential for CUORE and future tonne-scale cryogenic calorimeters as multipurpose experiments, demonstrating the scalability of this particle detector’s technology in view of both rare decays and dark matter interactions.